The people on thohoyandou Island buy only juice and cloth. The cip basket contains the quantities bought in 2010. The average household spent $60on the juice and $30 on cloth in 2009 when the price of juice wa $2 a bottle and the price of cloth was$5a meter. In current year 2010 juice is $4 a bottle and cloth is $6 a meter. Calculate cip per basket


1
Expert's answer
2021-09-08T08:06:51-0400

2009

basket value 

60+30=90 $

The CIP in 2009 = 90$/90$X100=100

60:2= 30 bottles of juice in a basket

30:5=6 cloth meter


2010

basket value = 30*4 + 6*6= 120 +36= 156 $


The CIP in 2010 = 156$/90$X100= 173


Now we can calculate the inflation rate between 2009 and 2010:

(173-100)/100=73/100=73%